OpenStudy (anonymous):

1 W!11 9!v3 4 M3d41 Indicate in standard form the equation of the line passing through the given point and having the given slope. D(5, -2) m = 2/5

OpenStudy (anonymous):

can ypu help?

OpenStudy (anonymous):

y+2=2/5(x-5)=2/5x -2 2/5x-y=4 would be a standart form

OpenStudy (anonymous):

it cant i have to fill in the blanks its like this _x__y___

OpenStudy (anonymous):

2/5x-1y=4

OpenStudy (anonymous):

2/5 doesnt fit it allows 1 input there

OpenStudy (anonymous):

2/5=0.4

OpenStudy (anonymous):

only 1 number fits before x and 2 numbers fit before y

OpenStudy (anonymous):

try: 2x-5y=20

OpenStudy (anonymous):

yes it was right thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line through (2, -1) and parallel to a line with slope of

OpenStudy (anonymous):

close this question and open new one, plz